# GSPRec: Temporal-Aware Graph Spectral Filtering for Recommendation

### v1

@@ B.1 GSPRec Algorithm
$\mathbf{S}^{\prime}\leftarrow\text{Symmetrize}(\mathbf{S})$ // $S^{\prime}_{ij}=1$ if $S_{ij}=1\text{ or }S_{ji}=1$
@@
$[\mathbf{U},\mathbf{\Lambda}]\leftarrow$ Eigendecomposition of $\mathbf{L}$
@@
$\bar{\boldsymbol{\lambda}}\leftarrow$ Diagonal entries of $\mathbf{\Lambda}$
$\mathbf{G}_{\text{BP}}\leftarrow\text{diag}\left(\exp\left(-\frac{(\bar{\boldsymbol{\lambda}}-c)^{2}}{w}\right)\right)$ // Gaussian band-pass
@@ B.2 Proof of Proposition 2: Laplacian Validity
Theorem 5 (Properties of Normalized Laplacian (Chung, 1997 [6])).
@@ Appendix C Theoretical Analysis of Spectral Bands
Mid-frequency components represent a balance where signals vary moderately across local graph neighborhoods - precisely the pattern exhibited by personalized preferences that differ between user communities but show consistency within them. This aligns with previous findings in spectral clustering [47] and network community detection [31], where mid-frequency eigenvectors have been shown to encode community structure.
We can formalize this intuition through the concept of total variation of a signal [43]. Given a signal $\mathbf{f}$ on a graph with Laplacian $\mathbf{L}$, the quadratic form $\mathbf{f}^{T}\mathbf{L}\mathbf{f}$ measures how smoothly the signal varies across the graph. For the eigenvectors $\mathbf{u}_{i}$ of $\mathbf{L}$, we have $\mathbf{u}_{i}^{T}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$, where $\lambda_{i}$ is the corresponding eigenvalue.
@@
In our empirical analysis across datasets, we consistently observe that coefficients $\alpha_{i}$ corresponding to mid-frequency components ($0.80.3<\lambda_{i}<$ 0.8) have significantly higher magnitude for user-specific preference patterns compared to global trends. This provides empirical validation of our theoretical understanding of spectral representation in recommendation graphs.
@@ Appendix D Computational Complexity Analysis
We analyze the computational complexity of GSPRec with $m$ users, $n$ items, $|\mathcal{D}|$ interactions, diffusion depth $d$, and $r$ eigenvectors.
D.1 Time Complexity
The time complexity of key components is:
  • Graph Construction: $O(|\mathcal{D}|+d\cdot\text{nnz}(\mathbf{S}^{\prime})\cdot n)$ for building and diffusing the sequential transition matrix.
  • Eigendecomposition: $O(r\cdot(m+n)^{2})$ for computing the $r$ smallest eigenvectors of the Laplacian.
  • Spectral Filtering: $O(|\mathcal{D}|\cdot r+(m+n)\cdot r^{2})$ for applying band-pass and low-pass filters.
The overall time complexity is dominated by the eigendecomposition step: $O(r\cdot(m+n)^{2})$. For sparse interaction data and $r\ll(m+n)$, the algorithm remains computationally tractable.
D.2 Space Complexity
The space requirements are:
  • Input Data: $O(|\mathcal{D}|+\text{nnz}(\mathbf{S}^{(d)}))$ for storing user-item interactions and sequential transitions.
  • Spectral Representation: $O(r\cdot(m+n))$ for storing eigenvectors.
  • Output: $O(m\cdot n)$ for the filtered results and final predictions.
The overall space complexity is $O(|\mathcal{D}|+\text{nnz}(\mathbf{S}^{(d)})+r\cdot(m+n)+m\cdot n)$, which is manageable for sparse real-world recommendation datasets where $|\mathcal{D}|\ll m\cdot n$ and $\text{nnz}(\mathbf{S}^{(d)})\ll n^{2}$.
In practice, the algorithm can be further optimized through sparse matrix operations, approximate eigendecomposition methods, and parallelization of filtering operations.
@@ Appendix E Baseline Details
  • LightGCN [14] simplifies GCN for recommendation by removing feature transformation and non-linear activation, retaining only essential neighborhood aggregation.
  • LR-GCCF [4] introduces residual connections to graph convolutional networks for improving recommendation performance.
  • IMP-GCN [23] propagates information in user sub-graphs to reduce the impact of noise and alleviate the over-smoothing problem.
  • SimpleX [28] focuses on optimizing loss function and negative sampling strategies.
  • UltraGCN [29] approximates infinite-layer graph convolutions via constraint loss and flexible edge weight assignment.
@@
  • GF-CF [42] integrates linear and ideal low-pass filters to model user preferences.
  • PGSP [25] uses a mixed-frequency low-pass filter over personalized graph signals.
  • FaGSP [52] employs Cascaded and Parallel Filter Modules for user preference modeling.
  • HiGSP [53] adopts cluster-wise and globally-aware filters to recognize unique and general interaction patterns.
For comprehensive evaluation, we report results for LR-GCCF, IMP-GCN, SimpleX, UltraGCN, and GF-CF as presented in [53, 52], which used identical dataset preprocessing and evaluation protocols.
@@ F.2.1 Approximate Eigendecomposition
The dominant computational bottleneck can be addressed using the Nyström method [8] or randomized SVD approaches [11], reducing complexity from $O(r(m+n)^{2})$ to $O(r^{2}(m+n))$ (or better, depending on specific variant, e.g., linear in $(m+n)$ for matrix-vector products), with minimal performance impact.
@@ F.2.2 Graph Sparsification
For large-scale graphs, we can apply spectral sparsification techniques [44] that preserve the essential spectral properties while reducing edge density, accelerating both graph construction and eigendecomposition.
@@ Appendix G Mathematical Analysis of Bandpass Filtering
This satisfies the formal definition of a bandpass filter: it achieves maximum response when $\bar{\lambda}=c$ (which corresponds to $\lambda_{c}=c(\lambda_{\max}-\lambda_{\min})+\lambda_{\min}$ in the original eigenvalue domain) and attenuates components with eigenvalues far from $\lambda_{c}$. For our normalized Laplacian with original eigenvalues $\lambda\in[0,2]$ (so $\lambda_{\min}\approx 0,\lambda_{\max}\approx 2$), setting $c\in(0,1)$ (the center in the normalized $\bar{\lambda}\in[0,1]$ domain) ensures the filter passes a band of mid-range eigenvalues while attenuating both low and high extremes. When applied to user preference signals decomposed as $\mathbf{x}=\sum_{i=1}^{m+n}\alpha_{i}\mathbf{u}_{i}$, the filtered output $\mathbf{y}=\mathbf{U}\text{diag}(g(\boldsymbol{\Lambda}))\mathbf{U}^{T}\mathbf{x}$ selectively preserves components $\alpha_{i}\mathbf{u}_{i}$ where the eigenvalue $\lambda_{i}$ (corresponding to $\mathbf{u}_{i}$) results in a high $g(\lambda_{i})$ value.
Our empirical analysis across recommendation datasets reveals that coefficients $\alpha_{i}$ corresponding to intermediate eigenvalues ($0.80.3<\lambda_{i}<$ 0.8, assuming $\lambda_{i}$ are original eigenvalues in $[0,2]$) have higher magnitude for user-specific preference patterns compared to global trends. This confirms that mid-frequency components encode valuable personalization signals that traditional low-pass approaches would suppress.

### v2

@@ B.1 GSPRec Algorithm
[…]
@@
$\mathbf{U},\mathbf{\Lambda}]\leftarrow$ Eigendecomposition of $\mathbf{L}$
@@
$\bar{\bm{\lambda}}\leftarrow$ Diagonal entries of $\mathbf{\Lambda}$
$\mathbf{G}_{\text{BP}}\leftarrow\text{diag}\left(\exp\left(-\frac{(\bar{\bm{\lambda}}-c)^{2}}{w}\right)\right)$ // Gaussian band-pass
@@ B.2 Proof of Proposition 2: Laplacian Validity
Theorem 5 (Properties of Normalized Laplacian (Chung, 1997 Chung (1997))).
@@ Appendix C Theoretical Analysis of Spectral Bands
Mid-frequency components represent a balance where signals vary moderately across local graph neighborhoods - precisely the pattern exhibited by personalized preferences that differ between user communities but show consistency within them. This aligns with previous findings in spectral clustering Von Luxburg (2007) and network community detection Newman (2006), where mid-frequency eigenvectors have been shown to encode community structure.
We can formalize this intuition through the concept of total variation of a signal Shuman et al. (2013). Given a signal $\mathbf{f}$ on a graph with Laplacian $\mathbf{L}$, the quadratic form $\mathbf{f}^{T}\mathbf{L}\mathbf{f}$ measures how smoothly the signal varies across the graph. For the eigenvectors $\mathbf{u}_{i}$ of $\mathbf{L}$, we have $\mathbf{u}_{i}^{T}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$, where $\lambda_{i}$ is the corresponding eigenvalue.
@@
In our empirical analysis across datasets, we consistently observe that coefficients $\alpha_{i}$ corresponding to mid-frequency components ($0.3<\lambda_{i}<0.8$) have significantly higher magnitude for user-specific preference patterns compared to global trends. This provides empirical validation of our theoretical understanding of spectral representation in recommendation graphs.
@@ Appendix D Computational Complexity Analysis
GSPRec has time complexity dominated by eigendecomposition and filtering operations. For sparse datasets with $r\ll(m+n)$, the method remains tractable on benchmark datasets as demonstrated by our runtime analysis (Table 7). For large-scale deployments, approximation techniques like polynomial filtering can reduce computational requirements to $O(K\cdot|\mathcal{D}|)$ for $K$-order polynomials, making the approach practical for industrial applications.
@@ Appendix E Baseline Details
  • LightGCN He et al. (2020) simplifies GCN for recommendation by removing feature transformation and non-linear activation, retaining only essential neighborhood aggregation.
  • LR-GCCF Chen et al. (2020b) introduces residual connections to graph convolutional networks for improving recommendation performance.
  • IMP-GCN Liu et al. (2021) propagates information in user sub-graphs to reduce the impact of noise and alleviate the over-smoothing problem.
  • SimpleX Mao et al. (2021a) focuses on optimizing loss function and negative sampling strategies.
  • UltraGCN Mao et al. (2021b) approximates infinite-layer graph convolutions via constraint loss and flexible edge weight assignment.
@@
  • GF-CF Shen et al. (2021) integrates linear and ideal low-pass filters to model user preferences.
  • JGCF Guo et al. (2023) uses Jacobi polynomial-based filters to amplify high-frequency signals that are typically suppressed by standard GCNs.
  • PGSP Liu et al. (2023) uses a mixed-frequency low-pass filter over personalized graph signals.
  • HiGSP Xia et al. (2024) adopts cluster-wise and globally-aware filters to recognize unique and general interaction patterns.
  • FaGSP Xia et al. (2025) employs Cascaded and Parallel Filter Modules for user preference modeling.
For comprehensive evaluation, We use official repositories for methods for reproducing results on our dataset settings. We report results for LR-GCCF, IMP-GCN, SimpleX, UltraGCN, and GF-CF as presented in Xia et al. (2024; 2025), which used identical dataset preprocessing and evaluation protocols.
E.1 Filter Design
@@ F.2.1 Approximate Eigendecomposition
The dominant computational bottleneck can be addressed using the Nyström method Fowlkes et al. (2004) or randomized SVD approaches Halko et al. (2011), reducing complexity from $O(r(m+n)^{2})$ to $O(r^{2}(m+n))$ (or better, depending on specific variant, e.g., linear in $(m+n)$ for matrix-vector products), with minimal performance impact.
@@ F.2.2 Graph Sparsification
For large-scale graphs, we can apply spectral sparsification techniques Spielman & Teng (2011) that preserve the essential spectral properties while reducing edge density, accelerating both graph construction and eigendecomposition.
@@ Appendix G Mathematical Analysis of Bandpass Filtering
This satisfies the formal definition of a bandpass filter: it achieves maximum response when $\bar{\lambda}=c$ (which corresponds to $\lambda_{c}=c(\lambda_{\max}-\lambda_{\min})+\lambda_{\min}$ in the original eigenvalue domain) and attenuates components with eigenvalues far from $\lambda_{c}$. For our normalized Laplacian with original eigenvalues $\lambda\in[0,2]$ (so $\lambda_{\min}\approx 0,\lambda_{\max}\approx 2$), setting $c\in(0,1)$ (the center in the normalized $\bar{\lambda}\in[0,1]$ domain) ensures the filter passes a band of mid-range eigenvalues while attenuating both low and high extremes. When applied to user preference signals decomposed as $\mathbf{x}=\sum_{i=1}^{m+n}\alpha_{i}\mathbf{u}_{i}$, the filtered output $\mathbf{y}=\mathbf{U}\text{diag}(g(\bm{\Lambda}))\mathbf{U}^{T}\mathbf{x}$ selectively preserves components $\alpha_{i}\mathbf{u}_{i}$ where the eigenvalue $\lambda_{i}$ (corresponding to $\mathbf{u}_{i}$) results in a high $g(\lambda_{i})$ value.
Our empirical analysis across recommendation datasets reveals that coefficients $\alpha_{i}$ corresponding to intermediate eigenvalues ($0.3<\lambda_{i}<0.8$, assuming $\lambda_{i}$ are original eigenvalues in $[0,2]$) have higher magnitude for user-specific preference patterns compared to global trends. This confirms that mid-frequency components encode valuable personalization signals that traditional low-pass approaches would suppress.